OpenStudy (anonymous):

Let f(x) = x2 - 3x - 7. Find f(-3). 11 8 -7 -10

OpenStudy (anonymous):

(-3)^2-3(-3)-7=9+9-7=11

OpenStudy (anonymous):

Thank you!!!!

OpenStudy (anonymous):

11